# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

### v1

@@ 4.3. Bootstrapping Sampling Strategy
4.4.2. Hyperparameter Tuning.
In our research, the primary focus is on the hyperparameter tuning of various collaborative filtering (CF) recommendation models, selected as baselines. These models are specifically tailored and evaluated within implicit feedback scenarios, where the subtleties of user preferences are inferred from indirect interactions. This approach ensures a comprehensive exploration of hyperparameter spaces to optimize model performance in these nuanced environments. The RecBole public library [^2] https://www.recbole.io/ is used for implementing and applying these models. Hyperparameter tuning is an essential step in optimizing the performance of these models. We employ a greedy search strategy to identify the best configurations based on their performance on the validation set. The following is a summary of the hyperparameter tuning conducted for each model:
@@
The total number of hyper-parameters is 14, encompassing a total of 223 experimental cases. This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our resear
5. Results
In this section, we present insights gained from the experimental validation of our system. We have divided our analysis into four key areas, each focusing on different aspects of performance: (i) personalization (cf. Section 5.1), (iii) stability of recommendations over different runs (cf. Section 5.2), (ii) provider fairness and diversity of recommendations (cf. Section 5.3), and finally (iv) the temporal freshness (or recency) of recommendations (cf. Section 5.4). We also provide an interesting discussion on the genre dominance/bias of recommendation models employed in Section 5.5. We approach this section by formulating a number of the hypotheses for each part, afterward showcasing the results, and providing the associated analysis. Finally, we will provide answers to these hypotheses.
5.1. Personalization of Recommendations
@@
This section focuses on the personalization performance measured by recommendation top-$K$ accuracy. Results can be found in Table 5, in which each generative scenario (i.e., using ChatGPT) could be found on the upper tabs while the result of CF baselines is on the bottom. It should be noted that, given the extensive nature of these experiments, not every possible scenario combination was tested independently; rather, a select subset of scenarios was examined.
We will analyze the results by evaluating the average performance (abbreviated as Avg Perf.) in terms of NDCG and Recall across different conditions/models. These conditions include exploring the impact of (i) the system’s role: No Role (R0), Standard Recommender (R1), and Fairness-aware Recommender (R2), and (ii) the presence (E1) or absence (E0) of a fairness statement in the LLM prompts.
Table 5. Table of the performance result measured in terms of recommendation Top-10 accuracy.
“System” Role: No (R0), Emphasis: No (E0)
Avg Perf. Bootstrap Mean Bootstrap Conf
@@
S7. Chain-of-thought (COT) 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
“System” Role: No (R0), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
@@
S7. Chain-of-thought (COT) 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
“System” Role: Normal Recommender (R1), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Conf
@@
S7. Chain-of-thought (COT) 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
“System” Role: Fair Recommender (R2), Emphasis: No (E0)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
@@
Additionally, a comparison between the regular performance values (obtained using average across values) and bootstrap averages of model performances provides deeper insights. The bootstrap method, which involves repeatedly resampling with replacement, offers a more robust understanding of the variability in performance measures. Table 5 illustrates this comparison. For instance, while the regular average provides a single estimate of NDCG and Recall, the bootstrap mean extends this by offering confidence intervals. These confidence intervals, represented as Bootstrap Conf. in the table, indicate the range within which the true performance measure is likely to fall with a certain level of confidence.
We outline below the hypotheses regarding personalization that are central to this section:
Hypothesis 1 (H1): The incorporation of a variety of goal-oriented prompts (as detailed in scenarios S1 to S7 in Section 3.1) across diverse settings (e.g., involving different system assignment roles) is anticipated to yield substantially varied performance outcomes in RecLLM. Hypothesis 2 (H2): GPT-based recommenders, particularly in a zero-shot setting, could achieve results comparable to CF baselines in specific scenarios. This is more probable in situations where the system is encouraged to engage in motivated reasoning and/or adopt COT strategies. This hypothesis is founded on observations that such strategies have demonstrated competitive performance in previous ML tasks.
The Impact of Prompt Structure and Setting
In the uppermost tab of Table 5, which presents the settings designated as No Role (R0) for the system role, the results are typically less favorable. Specifically, the prompt scenarios Simple (S1) and Motivated Reasoning (S5) demonstrate better performance within their respective segments. Scenario S1 under No Role (R0) reports NDCG and Recall values of 0.008803 and 0.00994, respectively. In comparison, S5 shows slightly improved metrics, with an NDCG of 0.010152 and a Recall of 0.010906. Overall, these results are the poorest when compared to other settings utilizing LLM-promoted prompts.
Contrastingly, the data in tabs 3 and 4 indicate that when roles are delineated as either Normal Recommender (R1) or Fair Recommender (R2), there is a notable enhancement in both NDCG and Recall metrics. It is of particular note that under these roles (R1 and R2), scenarios such as simple (S1), chain-of-thought (COT) (S7), and Motivated Reasoning (S6) markedly exceed the performance metrics of their No Role (R0) counterparts. The most pronounced performance increase is evident in the fourth tab, where models reach their optimal enhancements. The simple model (S1) advances from 0.008803 to 0.014268 in NDCG (approximately a 62% increase) and from 0.00994 to 0.015796 in Recall (approximately a 59% increase). The Motivated Reasoning (S6) and COT (S7) models also exhibit substantial gains, with S6 escalating from 0.010152 to 0.012354 in NDCG (21.7% increase) and from 0.010906 to 0.011968 in Recall (9.7% increase), and S7 rising from 0.006573 to 0.013212 in NDCG (101% increase) and from 0.005848 to 0.012862 in Recall (120% increase). This trend is similarly observed in the Normal Recommender, Emphasis: Fair (R2/E1) configuration. These results emphasize the significant advantages of integrating system roles into recommendation systems.
Bootstrap values in the R2-E0 setup further affirm the consistency and reliability of these improvements. Both the ‘simple’ and ‘COT’ models exhibit stable performance ranges across different roles and conditions.
Comparison with CF Baselines
When evaluating the effectiveness of GPT-based recommendation models compared to CF baselines, it is apparent that both the NDCG and the recall values for the GPT models are generally lower than those for the CF baselines. This suggests that, while GPT-based models have made significant strides, especially when designated with specific system roles such as “Normal Recommender” (R1) or “Fair Recommender” (R2), they still lag behind CF models in terms of both ranking relevant items at the top (NDCG) and retrieving a broader set of relevant items.
For example, even with the “Fair Recommender” (R2) role applied, the “Simple” (S1) GPT-based model achieves an NDCG of 0.014268 and a recall of 0.015796, which are enhancements of the scenario “No Role” (R0) but still do not surpass the CF baseline models such as “BPR-MF,” “ItemKNN,” “NCF,” and “LightGCN,” that demonstrate higher efficacy in both metrics. The “TopPop” model, while having the lowest scores among the CF baselines, still presents an NDCG and Recall that are competitive with the improved GPT-based models.
In summary, some noteworthy best and worst performances can be found below:
  • For the GPT-based models under the “Fair Recommender” (R2) role with no fairness emphasis (E0):
    • The Simple (S1) scenario achieves the best NDCG of 0.014268 and a Recall of 0.015796.
    • The Diversify (S4) scenario shows the worst NDCG of 0.002844 and a Recall of 0.004901.
  • For the CF baselines:
    • The ItemKNN model has the highest performance with an NDCG of 0.04905 and a Recall of 0.06558.
    • The TopPop model displays the lowest performance among the CF baselines with an NDCG of 0.030556 and a Recall of 0.032327.
Despite the clear impact of prompts on the performance of RecLLM, it can be noted that Top-$k$ recommendation performance is not on par with traditional CF baselines. There are multiple strategies to address this issue, including enhancing prompt clarity and transitioning to few-shot learning. However, we posit that a significant factor contributing to this discrepancy is the possibility that many movies recommended by the GPT-recommender might not be present in the dataset. Therefore, by constraining GPT to select items only from the test item set, we believe that this could markedly improve the accuracy and relevance of the recommendations.
H1: Supported. The introduction of specific roles (e.g., Normal and Fair Recommender) in recommendation systems significantly enhances the performance of generative models. This is evidenced by substantial increases in NDCG and Recall metrics when comparing the No Role scenario (R0) with Role-Based scenarios (R1, R2). It is worth noting in terms of which prompt scenario is effective, the Simple (S1) scenario, suggesting “Recommend 10 movies that the user will likely enjoy,” appears to provide the most effective approach for achieving the best top-$K$ recommendation accuracy. In other scenarios, only the reasoning scenario, such as COT (S7) matches this performance. (H2): Partially Supported. The evidence does not support Hypothesis 2 (H2), which posited that GPT-based recommenders could match CF baselines in certain settings. While GPT models, particularly with system roles and strategies scuh as Simple, ‘Motivated Reasoning’ and ‘COT’, show enhanced NDCG, Recall, they generally do not reach the performance of CF baselines. It is worthwhile that these results indicate a trend towards achieving outcomes closer to those of Top-Pop with the refinement of prompt design, better rationalizing, or transitioning to a few-shot learning setting.
5.2. Stability and Confidence of Personalization Metrics
@@
As stated in Section 3.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models. To study the impact of such randomness, each experiment was conducted five times specifically for GPT-based models. It is important to note that prior to our experiments, we lacked specific knowledge about the performance of Chat-GPT prompts. Therefore, our experiments focused exclusively on the R0/E0 scenario, which demonstrated poorer performance compared to the CF baseline. Despite this, the variability of results across different runs is noticeable. Consequently, in the analysis of this section, we should focus on observing the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
Hypotheses on Stability of the Top-$K$ Accuracy Metric
@@
Hypothesis 1 (H1): Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent NDCG scores.
@@
  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std333standard deviation of approximately 0.000305.
@@
While the current study demonstrates a high degree of stability in model results over multiple runs, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. The dynamic nature of data and algorithms in generative models means that updates and changes over time could lead to different behaviors and outputs. Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the model’s performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
@@
H1: Supported. The data presented in the NDCG graph aligns with the hypothesis. While there is some expected variability across multiple runs, it remains within an acceptable/predictable range. Confidence intervals overlap significantly for different runs of each GPT-based model. While this consistency across runs confirms the system’s reliability in providing personalized recommendations, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. Therefore, it is advisable to approach these results with caution when considering long-term deployment.
5.3. Diversity and Provider Fairness
This section examines the diversity and fairness of different recommender models. Fairness is assessed by measuring the equitable distribution of recommendations across all items, avoiding bias towards popular or specific subsets (Gini Index, HHI). Diversity is gauged through the variety and uniqueness of the recommendations, ensuring the inclusion of a broad spectrum of items (Entropy), refer to Section 4.2. These evaluations are crucial to understanding each model’s effectiveness in matching user preferences and its capacity to foster a balanced and varied recommendation environment.
H1: Given their training on extensive internet data, GPT-based models are hypothesized to exhibit a greater popularity bias than CF models, potentially leading to lower diversity in recommendations. This will be assessed using the Gini Coefficient and HHI for fairness, and Entropy for diversity. H2: The inclusion of a Fair Recommender as a ”system” role in GPT-based systems is expected to mitigate popularity bias and enhance the diversity of the system’s output.
For H1, upon examining the results presented, in Table 6 a notable observation is that the ItemKNN model consistently exhibits the highest recommendation fairness (achieving the lowest Gini Coefficient and highest Entropy) among all models. This, to conduct a focused discussion, we focused on the remaining machine learning-based CF models and GPT-based models, we will exclude ItemKNN from subsequent comparisons.
Without considering ItemKNN, the general trend suggests that the top-performing GPT models (Simple, Motivate Reasoning, and Chain-of-thought (COT)) are somewhat comparable to the majority of ML-based CF models, except when compared to NGCF, which shows a notably better performance in terms of fairness (Gini Coefficient: 0.950845 vs. 0.0982346, Entropy: 6.420996 vs. 5.042821) and diversity. Intriguingly, methods designed to promote diversification and novelty in GPT case (Surprise and Diversify Recommendation) perform contrarily, evidenced by their increased Gini Coefficients (Surprise: 0.997906, Diversify Recommendation: 0.992349) and reduced Entropy (Surprise: 3.227737, Diversify Recommendation: 4.232139), which is almost the worst in the respective scenarios, suggesting that these strategies may not effectively align with the objectives of reducing bias and enhancing diversity.
Therefore, upon excluding the exceptional performance of ItemKNN, which seems to leverage its distinct mechanism effectively in a dataset characterized by popularity bias, the GPT-based models’ levels of fairness and diversity appear comparable to those of CF models. The data suggests that GPT-based models generally match their CF counterparts, with exceptions such as NGCF which shows superior fairness. Interestingly, NGCF, despite being the least performing CF model in some aspects, still outperforms GPT-based models on fairness. When we turn our attention to the top-performing CF models, namely BPR and LightGCN, it is evident that they achieve their enhanced performance by concentrating user preferences on a narrower selection of items, thus limiting diversity—a trend less likely to be observed in GPT models. This is attributable to the GPT models’ design, which is not inherently influenced by the popularity bias present in our datasets.
For H2, the adaptation of GPT-based models with a ‘Fair Recommender’ system role results in a significant improvement in fairness and diversity metrics. Excluding ItemKNN for its good performance, the top GPT models exhibit better fairness in the Fair Recommender scenario, as evidenced by a noticeable decrease in Gini Coefficient (Simple: from 0.982463 to 0.978925, Chain-of-thought (COT): from 0.986889 to 0.979313) and HHI, coupled with an increase in Entropy (Simple: from 5.042821 to 5.387465, Chain-of-thought (COT): from 4.619500 to 5.365294). This trend suggests that GPT-based models, when guided by the system role within the ChatGPT framework, demonstrate a promising capacity for controllability. Such adaptability could mitigate potential biases inherent in the models—a flexibility that may not be as pronounced in traditional CF models.
However, specific strategies such as Diversification (S4) and Surprise (S5) integrated into GPT-based models do not uniformly follow this positive trend. For example, in the Diversify scenario, we observe a slight increase in the model’s Gini coefficient (from 0.992349 to 0.992603) and a marginal rise in HHI (from 0.034724 to 0.030010), which deviates from the expected reduction. Similarly, the Surprise scenario shows a rise in Gini Coefficient (from 0.997906 to 0.998365) and HHI (from 0.059857 to 0.067952), alongside a reduction in Entropy (from 3.227737 to 3.023948). These anomalies highlight that certain approaches within the Fair Recommender role may inadvertently intensify the biases they aim to alleviate.
The broader implication of this study is that the system role in the ChatGPT framework, alongside certain prompt scenarios, has the potential to align or even surpass the fairness and diversity benchmarks set by classical CF models. In notable instances, the bias is reduced, as indicated by lower Gini and HHI values, and increased Entropy. This indicates a potentially significant step forward in the controllability of GPT-based models to prevent their biases, an attribute that arguably gives them an edge over CF models in managing the diversity and fairness of recommendations.
Answer to H1: Partially supported. The results do not uniformly support H1. While certain GPT-based models do exhibit higher popularity bias as indicated by the Gini Coefficient and HHI compared to some CF models, this is not a consistent finding across all models. For example, models such as NGCF show superior fairness indicators compared to some GPT-based models. However, when excluding the outlier performance of ItemKNN, GPT-based models are generally on par with CF models regarding fairness and diversity. Therefore, the hypothesis that GPT-based models inherently lead to lower diversity due to popularity bias is only partially supported. Answer to H2: Supported. The evidence supports H2 to a significant extent. When GPT-based models adopt a Fair Recommender role, there is a noticeable improvement in fairness and diversity metrics, such as a decrease in Gini Coefficient and HHI, along with an increase in Entropy. This indicates that GPT-based models have an inherent flexibility that, when leveraged with the right system role, can effectively mitigate popularity bias and enhance the diversity of recommendations. In both instances, it is critical to acknowledge that specific strategies, namely “diversification” and “surprise,” do not conform to their anticipated effectiveness—in some cases, they produce outcomes contrary to expectations. This observation implies that certain types of prompts may inadvertently introduce complexity, leading to a confounding of the system’s performance.
Table 6. Diversity and provider fairness of the experimented models

### v3

@@ 4.1.3. Bootstrapping Sampling Strategy
Hyperparameter Tuning.
In our research, the primary focus is on the hyperparameter tuning of various collaborative filtering (CF) recommendation models, selected as baselines. These models are specifically tailored and evaluated within implicit feedback scenarios, where the subtleties of user preferences are inferred from indirect interactions. This approach ensures a comprehensive exploration of hyperparameter spaces to optimize model performance in these nuanced environments. The RecBole public library [^3] https://www.recbole.io/ is used for implementing and applying these models. Hyperparameter tuning is an essential step in optimizing the performance of these models. We used a bootstrapping sampling strategy, generating 1000 samples by repeatedly resampling our dataset with replacement.
The following is a summary of the hyperparameter tuning conducted for each model:
@@
The total number of hyper-parameters is 14, encompassing a total of 223 experimental cases. This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our research.
4.2. Experiment 2. Sequential Recommendation
4.2.1. Datasets and Setup
In this study, our objective is to understand the generalizability of the findings from the previous study and address aspects that were not studied previously. In particular, Experiment 2 focuses on the more practical task of sequential recommendation and the comparison of zero-shot vs. few-shot in-context learning (ICL), which we evaluate on datasets of different domains (movies and music). The choice of sequential recommendation, rather than classical top-$k$, was motivated by the research carried out in (Hou et al., 2023). Therefore, the primary goal in Experiment 2 is not to assess the quality of different prompt scenarios (as carried out in Section 4.1), rather to evaluate the effectiveness of zero-shot vs. few-shot ICL on the accuracy vs. item fairness of RS.
We focus our attention on two main datasets from the movie and music domains: MovieLens-1M and LastFM-1K. These datasets are widely recognized in the academic community, and provide a basis for evaluating our models in terms of fairness across different types of media consumption data. Here, we opt to use a sub-sample of users instead of all, to speed up experiments. Initially, we randomly select a subset of 80 users who exhibit a moderate level of interaction within the datasets. This allows us to handle the data efficiently while ensuring that the users selected have enough interactions to inform the training process. The data for these users is divided into training and test sets by sorting their interactions over time and splitting them so that 80% of a user interactions are used for training, with the remaining 20% held out for testing. This method respects the chronological order of interactions, thereby simulating a realistic scenario where a model can only learn from past data to make predictions about future user behavior.
Table 6. Statistics of the final datasets used in the sequential ICL task.
Dataset —U— —I— —R— $\frac{R}{U}$ $\frac{R}{I}$ Density Item Gini User Gini
LastFM-1K 80 5,500 277,607 3,470.09 50.47 98.6% 0.697 0.621
MovieLens 1M 80 3,900 20,987 262.34 5.38 99.2% 0.745 0.657
4.2.2. Baselines.
In this experiment, we do not compare the built RecLLMs with traditional sequential models, as this would require a separate and labor-intensive set of experiments, which is beyond the scope of this study. Instead, we compare the fairness and accuracy of different versions of RecLLMs against each other. Additionally, in one part of the experiment, we explore the impact of the presence or absence of user demographics in the prompts on recommendation fairness and quality. Since CF baselines typically do not use such demographic information, we exclude them from this comparison.
4.2.3. Evaluation metrics
The first experiment employs a diverse range of metrics to assess the performance of our recommendation system. These metrics are categorized into two primary classes: accuracy metrics, and item fairness, and catalogue coverage.
Accuracy Metrics.
For accuracy, we choose simpler metrics since the goal is not to compare them with strong CF baselines. We use the following two metrics to measure the quality of the top-$k$ recommendations in sequential setting:
  • Hit Rate (HR@k): Measures the proportion of users for whom at least one of the top-$k$ recommended items is relevant. It is calculated as:
[…]
    where $R_{Q}$ represents the set of recommendations of the ranker in question and $\mathcal{G}$ represents the set of items in the ground truth.
  • Average Rank (AverageRank@K): Measures the average position of relevant items in the recommendation list. It is calculated as:
[…]
    where $\text{rank}_{R_{Q}}(i)$ is the rank of item $i$ in the recommendation list $R_{Q}$.
Item fairness, catalouge Coverage and long-tail distribution.
Similar to the previous experiment, we use the Gini Index, HHI, and Entropy to measure item fairness. Additionally, we compute catalogue coverage, and display the long-tail distribution to visualize the concentration bias of recommendated items.
4.3. Parsing Output Recommendations and Finding Closest Match Items
Since language models (LLMs) naturally produce text-based outputs, we need a systematic approach to parse these outputs and map them to our existing dataset items. This ensures that when the RecLLM generates recommendations for which we do not find an exact match in the dataset, we can associate them with the most similar existing items, maintaining the integrity and relevance of our recommendation system.
  • We start by utilizing the OpenAI GPT-3.5 API itself to parse the raw text of the recommendations. We designed a LLM_parsing function to extract the recommended songs and artists (or movies) from the text provided by the language model. For instance, we specify a system message to format the output in a structured way, ensuring that each recommendation follows the format: “Song Name by Artist.” This structured output helps the subsequent matching process. The parsed recommendations are then collected, which contain cleaned and structured text.
  • Once the recommendations are parsed, the next step involves finding the closest matches to these recommendations within our dataset. This is achieved using the find_closest_match function, which employs the difflib.SequenceMatcher from the Python standard library. This tool calculates similarity scores between the recommended song and artist names and the existing items in our dataset. For each recommendation, the function compares it against all items in the dataset, computing a combined score based on the similarity of both the track name and the artist name.
  • A threshold is set to ensure that only matches with a high enough similarity score are considered valid. If the highest score exceeds this threshold, the corresponding item ID from the dataset is selected as the closest match. This method ensures robustness by preventing low-similarity matches from being accepted, thus maintaining the relevance and accuracy of the recommendations.
In summary, the implementation of these methods leverages essential Python tools and libraries, openai for API interaction, and difflib for sequence matching. These tools provide a robust and efficient framework for parsing and matching recommendations, ensuring that the RecLLM system can be effectively evaluated even when it generates items not originally present in the dataset.
5. Results and Discussion
In this section, we present the results and discussion for two experiments. We formulate hypotheses for each part, then showcase the results and provide analysis. Finally, we answer the hypotheses.
5.1. Experiment 1. Zero-Shot Top-k Recommendation
Through experiment 1, conducted in a zero-shot setting for the classical top-$k$ recommendation, we aim to answer the following experimental research questions. Details about Experiment 1 can be found in Section 3.1 and Section 4.1.
We evaluate the average performance (Avg Perf.) in terms of NDCG and Recall across different conditions/models. We compare regular performance values (average) with bootstrap averages to gain deeper insights. The bootstrap method involves repeated resampling and offers a more robust understanding of performance variability. Table 7 illustrates this comparison. While the regular average provides a single estimate, the bootstrap mean extends this by offering confidence intervals, indicating the range within which the true performance measure is likely to fall.
Experimental Research Questions RQ1. How does the incorporation of various goal-oriented prompts impact the accuracy (personalization) of GPT-based model in comparison to CF baselines? (cf. Section 5.1.1) RQ2. What is the stability and consistency of the personalization performance metrics (e.g., NDCG, Recall) for GPT-based recommendation systems across multiple runs? How does this variability compare to CF baselines? (cf. Section 5.1.2) RQ3. How do GPT-based recommendation systems compare to CF baselines in terms of provider fairness? Can the inclusion of a “Fair Recommender” system role might mitigate item unfairness and enhance diversity? (cf. Sections 5.1.3) RQ4. Do GPT-based recommendation systems exhibit a bias towards recommending newer or older movies, or certain genre tendancy, compared to CF baselines, and how does the temporal freshness of recommendations hold? How does the inclusion of explicit user ratings in prompts affect the personalization performance? (cf. Sections 5.1.4). RQ5. How do different algorithmic in-context learning (ICL) strategies, — in particular zero-shot vs. few-shot learning – impact the quality and biases of RecLLMs? What are the specific prompt design aspects that may influence this towards better or worse performance? (cf. Section 5.2) RQ6. What are the economic and practical implications of using GPT-based models for recommendation systems, specifically in terms of inference costs and latency issues? (cf. Section 5.3)
5.1.1. Personalization of Recommendations
@@
This section focuses on the personalization performance measured by recommendation top-$K$ accuracy. Results are presented in Table 7, with each generative scenario (using ChatGPT) in the upper sections and CF baselines in the bottom section. Not every possible scenario combination was tested; rather, a select subset was examined.
Hypotheses
To investigate the effectiveness of different recommendation strategies, we propose the following hypotheses:
  • H1. Incorporating goal-oriented prompts, including personalization-based, beyond-accuracy, and reasoning-based prompts (as detailed in scenarios S1 to S7 in Section 3.1.1), with different system assignment roles, results in substantially varied performance outcomes in RecLLMs;
  • H2. In terms of performance, GPT-based recommenders in zero-shot setting can provide comparable performance compared to collaborative filtering (CF) baselines.
Discussion.
Our evaluation highlights several findings regarding the personalization of recommendations using RecLLM.
  • First, there is a clear indication that incorporating the “system role” improves the performance of RecLLM, and this consistency is observed across most scenarios. For example, for a randomly chosen scenario S1, in Tab1 and Tab2 (with no system role assignment), the NDCG performances are 0.008803 and 0.010431, respectively. In contrast, Tab3 and Tab4 (with system role assignments, either as “act as a recommender system” or “act as a fair recommender system”) show performance improvements to 0.013007 and 0.014268, respectively. This represents roughly a 45% improvement. In some scenarios, the improvement reaches up to 100%. It becomes thus evident that the system role plays a crucial part in optimizing the effectiveness of recommendation systems.
  • Second, the performance varied considerably across different scenarios (S1 to S7), indicating that the context and nature of the prompts affect recommendation outcomes. For example, in Tab4, the “Simple” prompt scenario (S1) achieved an NDCG of 0.014, while the “Diversify” prompt (S4) had a lower NDCG of 0.002. The best-performing methods were simple, reasoning-based (COT and motivated reasoning), providing relatively better accuracy. Thus, based on these results, we may conclude that requesting ChatGPT to diversify recommendation or to provide novel recommendations significantly reduces accuracy, which should be avoided or carefully considered. Overall these variations underscore the importance of selecting appropriate prompts based on the desired recommendation attributes.
  • Finally, comparing RecLLMs with baselines, our results reveal that RecLLM, even in zero-shot settings, generally performs lower than CF baselines, though in certain contexts, this difference shrinks considerably. For instance, the average NDCG of CF models such as BPR-MF and ItemKNN were 0.04776 and 0.04905, respectively. In comparison, the best-performing RecLLM prompt achieved an NDCG of 0.014, which is approximately 70% lower.
Table 7. Performance results measured in terms of recommendation Top-10 accuracy.
Tab 1 - System Role: No (R0), Emphasis: No (E0)
Prompt: Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy.
System Role: -
Avg Perf. Bootstrap Mean Bootstrap Conf.
@@ Discussion.
S7. COT 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
Tab 2 - System Role: No (R0), Emphasis: Fair (E1)
Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend … .
System Role: -
Avg Perf. Avg. Bootstrap Bootstrap Conf.
@@
S7. COT 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
Tab 3 - System Role: Normal Recommender (R1), Emphasis: Fair (E1)
Prompt: Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy.
System Role: Act as a fair recommender system balancing between popular and less-known movies to ensure provider fairness.
Avg Perf. Avg. Bootstrap Bootstrap Conf.
@@
S7. COT 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
Tab 4 - System Role: Fair Recommender (R2), Emphasis: No (E0)
Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend … .
System Role: Given a user, act as recommender system.
Avg Perf. Avg. Bootstrap Bootstrap Conf.
@@
Answer to Hypotheses
To address the proposed hypotheses, we present the following findings:
  • H1. $\checkmark$ Supported – The integration of specific system roles (e.g., “act as a recommender,” or “act as fair recommender”) leads to considerable performance improvements in RecLLMs.
  • H2. $\times$ Rejected – Performance of RecLLMs in zero-shot settings consistently and significantly lags behind CF baselines.
5.1.2. Stability and Confidence of Personalization Metrics
@@
As stated in Section 3.1.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models. To study the impact of such randomness, each experiment was conducted five times specifically for GPT-based models. [^4] Note that prior to these experiments, we lacked specific knowledge about the performance of Chat-GPT prompts. Our experiments thus focused on a randomly selected case (R0/E0 scenario). In the analysis of this section, our goal is to observe the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
Hypotheses
To rigorously evaluate the robustness and consistency of GPT-based models under varying conditions, we propose the following hypothesis:
@@
  • H3. Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent recommendation quality.
Discussion
@@
  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std555standard deviation of approximately 0.000305.
@@
While the current study demonstrates a high degree of stability in model results over multiple runs, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. The dynamic nature of data and algorithms in generative models means that updates and changes over time could lead to different behaviors and outputs. Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
Answer to hypotheses
To evaluate the validity of our proposed hypotheses, we present detailed findings below, supported by statistical data and analyses:
@@
  • H3. $\checkmark$ Supported – This hypothesis, which anticipated consistent performance across different GPT-based model runs, is supported. The data presented in the NDCG graph aligns with this expectation, showing only minor variability across multiple runs. Confidence intervals overlap significantly for different runs of each GPT-based model, which confirms the reliability of the system in providing personalized recommendations. However, it is important to consider these results with caution for long-term deployment as they do not necessarily guarantee stability over extended periods such as weeks or months.
5.1.3. Item Fairness
This section explores the item fairness metrics of different recommender models, crucial for evaluating their effectiveness in providing balanced and varied recommendations.
Hypotheses
To ascertain the impact of different model strategies on recommendation fairness, we carefully analyzed the outcomes in the context of our formulated hypotheses:
  • H4: Due to their training on large-scale internet data, GPT-based models are hypothesized to exhibit lower item fairness than classical CF models. However, asking the system to produce more diverse or additional recommendations (scenarios S3 and S4) is expected to increase/improve item fairness.
  • H5: Integrating a “Fair Recommender” as a system role in GPT-based systems is expected to enhance item fairness.
Table 8. Provider fairness of the tested models in Experiment 1